OpenStudy (anonymous):

what is the intersection point of y-x = 0 and y-3x = 6 ?

OpenStudy (ajprincess):

y-x=0 y=x y-3x=6 Substitute y=x in the second equation x-3x=6 find x from this. then plug in the value of x in either of the equation to find y.

OpenStudy (ajprincess):

Does that help? @missylulu

OpenStudy (anonymous):

yesh, thanks so much @ajprincess !!!

OpenStudy (ajprincess):

Welcome:)